ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 152, 2006

Abraham, U., I. Gorelic and I. Juhász On Jakovlev spaces  205
Babson, E. and D. N. Kozlov Complexes of graph homomorphisms 285

Chatterjee, P. and D. W. Morris Divergent torus orbits in homogeneous spaces of ℚ-rank two

229

Cheung, Y. and H. Masur A divergent Teichmüller geodesic with uniquely ergodic vertical foliation

1

Cutillas Ripoll, P. On infinite products of functions in compact Riemann surfaces

349

Defant, A. and L. Frerick A logarithmic lower bound for multidimensional Bohr radii

17

Farb, B. and G. C. Hruska Commensurability invariants for nonuniform tree lattices

125

Ferenczi, V. and E. M. Galego Some equivalence relations which are Borel reducible to isomorphism between separable Banach spaces

61

Frerick, L. See A. Defant

Galego, E. M. See V. Ferenczi

Gorelic, I. See U. Abraham

Hruska, G. C. See B. Farb

Juhász, I. See U. Abraham

Konyagin, S. and I. Łaba Distance sets of well-distributed planar sets for polygonal norms

157

Koslov, D. N. See E. Babson

Krushkal, S. and R. Kühnau Grunsky inequalities and quasiconformal extension

49

Kühnau, R. See S. Krushkal

Kulagin, V. Classification of cocycles of generic equivalence relations  83

Łaba, I. See S. Konyagin

Marcus, M. and L. Véron Maximal solutions of semilinear elliptic equations with locally integrable forcing term

333

Masur, H. See Y. Cheung Maucourant, F. Dynamical Borel–Cantelli lemma for hyperbolic

spaces

143

Morris, D. W. See P. Chatterjee

Muchnik, R. A note on stationarity of spherical measures  271
Neeman, I. Unraveling Π_{1}^{1} sets, revisited  181

Rödl, V., M. Schacht, E. Tengan and N. Tokushige Density theorems and extremal hypergraph problems

371

Schacht, M. See V. Rödl

Shelah, S. and A. Usvyatsov Banach spaces and groups—Order properties and universal models

245

Silver, D. S. and S. G. Williams Lifting representations of

ℤ-groups 313

Sodin, M. and B. Tsirelson Random complex zeroes, II. Perturbed lattice

105

Tengan, E. See V. Rödl

Tokushige, N. See V. Rödl

Tsirelson, B. See M. Sodin

Usvyatsov, A. See S. Shelah

Véron, L. See M. Marcus

Weiss, B. Divergent trajectories and ℚ-rank  221

Williams, S. G. See D. S. Silver

Zapletal, J. Proper forcing and rectangular Ramsey theorems  29